OpenStudy (drummmlers):

The length of a room is 6 feet less than twice the width and the area of the floor is 260 square feet. What are the dimensions of the room?

OpenStudy (anonymous):

\[L=2w-6\]\[A=lw=260\]\[260=w(2w-6)\]\[2w^{2}-6w-260=0\]\[w^{2}-3w-130=0\]\[(w-13)(w+10)=0\]\[w=13\]\[l=20\] To check\[260=13\times20\]\[260=260\]CHECK